OpenStudy (anonymous):

Assuming the balls to be identical except for difference in colours, the number of ways in which one or more balls can be selected from 10 white, 9 green and 7 black balls is : (1) 880 (2) 629 (3) 630 (4) 879

OpenStudy (anonymous):

(10+1)*(9+1)*(7+1)-1= 11*10*8-1= 880-1 =879